Dinner guests ate 25 percent of an apple pie and 60 percent of a cherry pie. In all, how much pie was left?
Answer by checkley77(12844)

1-.25=.75 OR 75% OF THE APPLE PIE WAS LEFT.
1-.60 .40 OR 40% OF THE CHERRY PIE WAS LEFT.
OR:
(.75+.40)/2=1.15/2=57.5% OF THE 2 PIES WERE LEFT.